### v3

# spacing=nonfrench

Abstract.
5G mmWave roadside networks promise high bandwidth but face significant challenges in maintaining reliable connections for users moving at high speed. Frequent handovers, complex beam alignment, and signal blockage from car bodies lead to service interruptions and degraded performance. We present Wall-Street, a vehicle-mounted smart surface that enhances 5G mmWave connectivity for in-vehicle users. Wall-Street improves mobility management by (1) steering outdoor mmWave signals into the vehicle for shared coverage, enabling a single, collective handover decision for all users; (2) enabling neighbor-cell search without interrupting data transfer, allowing for seamless handovers; and (3) connecting users to a new cell before disconnecting from the old cell for reliable handovers. We have implemented and integrated Wall-Street in the COSMOS testbed and evaluated its real-time performance with multiple gNBs and users inside a surface-mounted vehicle, driving on a nearby road. In multi-UE scenarios, Wall-Street improves TCP throughput by up to 78% and reduces RTT by up to 34% over the standard 5G Standalone handover protocol.
1. Introduction
5G wireless networks leverages millimeter wave (mmWave) spectrum in the FR2 band to deliver multi-gigabit, low-latency connectivity (3GPP, 2022). However, due to its short wavelengths and limited diffraction capability, mmWave spectrum suffers from high path loss and extreme vulnerability to blockage, posing significant challenges in maintaining reliable connections, particularly in high mobility scenarios (Rappaport et al., 2017).
To overcome poor signal propagation, mmWave networks require dense deployments of small cells, which in turn leads to handovers (HOs) that are far more frequent than in low-band networks. Indeed, mmWave 5G NR experiences a HO significantly more often than LTE or low band 5G NR (Hassan et al., 2022; Narayanan et al., 2021). The mmWave HO process adds a layer of complexity on top of low-band procedures (Deng et al., 2018; Xu et al., 2020), requiring precise alignment of narrow beams between the base station (gNB) and user equipment (UE) (Tassi et al., 2017). To achieve this, the standard HO protocol requires the UE to frequently suspend its data connection to measure neighboring gNBs, a procedure that directly leads to service interruptions and performance degradation.
High delays associated with mmWave HOs have significant impacts throughout the entire protocol stack, from the physical layer to the application layer, degrading quality of service (QoS) for UEs, especially for latency-sensitive applications (Hassan et al., 2022; Narayanan et al., 2020). To prepare a HO, the UE immediately stops data communication with the serving gNB to perform neighboring cell measurements. This process requires UEs to engage in an exhaustive search for the strongest signal from neighboring gNBs, resulting in bufferbloat and slow growth of the congestion window (Mezzavilla et al., 2018). While existing works (Wang et al., 2020) primarily focus on reducing beam searching overhead, the fundamental problem of maintaining reliable connectivity remains unexplored. Also, the HO decision relies on these measurement from UEs, which may be outdated or lost due to the rapidly fluctuating channel conditions. This can trigger flawed decisions, leading to ping-pong handovers and connection timeouts, impacting overall network performance.
This paper presents the design and implementation of Wall-Street, a programmable smart surface strategically deployed on a vehicle’s interior to address these challenges. Wall-Street acts as an intelligent gateway for mmWave signals, overcoming vehicle body blockage and redesigning the HO process to eliminate service interruptions for in-vehicle users. It achieves this through a new capability called dual-beam control; Wall-Street steers two incoming beams simultaneously in reflective and/or transmissive directions, with arbitrary power division between them. Using this ability, it offloads the HO task from individual UEs to the surface itself. This allows all in-vehicle users to maintain data connectivity while the network performs HO process on their behalf. As illustrated in Fig.˜1, we summarize three key innovations:
Without Wall-Street:     With Wall-Street:
(a) Enhance in-vehicle mmWave penetration.
(b) Transflexive handover, ongoing comm.
(c) Enhance handover reliability.
Figure 1. Wall-Street’s design innovations electronically split, shape, and steer mmWave transmissions in real time to enable seamless roadside 5G networks.
1. Batched handover via shared in-vehicle coverage. Wall-Street provides consistent, shared mmWave coverage to all UEs by steering outdoor signals into the vehicle (Fig.˜1(a)). This also allows the network to batch the HO procedure for all in-vehicle UEs into a single operation, eliminating the overhead and contention of separate HOs.
2. Seamless handover with concurrent communication. Innovating beyond existing surfaces that can only reflect or refract at one time (Cho et al., 2023), Wall-Street uniquely steers two mmWave beams simultaneously in transmissive and reflective angles with minimal signal loss. Specifically, it reflects neighboring gNB’s beacon signals to the serving gNB for measurement while keeping the user data link through the transmitted path [^1] To avoid collisions, nearby gNBs separate their respective signals in frequency and time across different gNBs, allowing the source gNB to listen to other cells while simultaneously exchanging data with the UE. (Fig.˜1(b)). This eliminates the standard protocol’s requirement to suspend data transmission during cell search, removing service interruption. Since users are not involved in the scanning process, it also minimizes UE energy consumption related to HOs, which typically consumes twice the energy compared to low-band HOs (Hassan et al., 2022).
3. Make-before-break handover. At the moment of HO, Wall-Street combines beams from both the source and target gNBs. By steering duplicate data packets from both gNBs to the user, it significantly reduces packet loss and retransmissions, enhancing reliability during the transition. Also, by adjusting the power of the respective links, Wall-Street ensures an uninterrupted flow of data.
We have designed, implemented, and integrated the full Wall-Street system, including its hardware and protocol, into the COSMOS wireless testbed (Raychaudhuri et al., 2020). We have also implemented new PHY features on the testbed for HO operations. We mount Wall-Street on the rear door of an SUV with multiple mobile UEs inside, communicating with three gNBs placed on the first floor of a lab facing the road. To evaluate TCP end-to-end performance, we use a state-of-the-art mmWave ns-3 simulator  (Mezzavilla et al., 2018) with our own modifications. The simulation is trace-driven, fed with PHY traces collected from real-time experiments. We compare our system against a standard 5G Standalone HO protocol that we also implemented as a baseline. In multi-user scenarios, our results demonstrate that Wall-Street improves TCP throughput by up to 78%, reduces RTT by up to 34%, and cuts unnecessary HOs in half. Microbenchmarks validate Wall-Street’s in-vehicle coverage improvement, beam tracking, and dual-beam steering.
Contribution. Our main contributions are as follows:
  • $\bullet$
    A dual-link beam manipulation. We design, build, and demonstrate the first mmWave smart surface capable of simultaneously and independently steering a transmissive link (into the vehicle) and a reflective link (between gNBs) with programmable power division.
  • $\bullet$
    Reliable make-before-break handover. We realize a make-before-break HO protocol that leverages our surface to connect UEs to a new gNB before disconnecting from the old gNB. By receiving packets from both gNBs, our system improves HO reliability.
  • $\bullet$
    Scalable, vehicle-centric handover. We introduce an algorithm that makes a single, batched HO decision for all in-vehicle users, reducing signaling overhead. By estimating the stable gNB-to-surface link quality from multi-user measurements, our design also improves HO decision.
  • $\bullet$
    Real-world system implementation. We implement the full Wall-Street hardware system and integrate it in the COSMOS testbed. Our experiments with a surface-mounted SUV and multi-gNBs/UEs reveals significant improvement over the standard protocol.
2. Related Work
High-speed Mobility Management. Several works address sub-6 GHz vehicular mobility. Wi-Fi Goes to Town (Song et al., 2017) provides continuous connectivity via rapid switching between roadside Wi-Fi APs. POLYCORN (Ni et al., 2023) optimizes traffic scheduling for high-speed railways but requires fixed trajectories, making it unsuitable for unpredictable vehicular movement. REM (Li et al., 2020) uses multipath profiles instead of signal strength for HO decisions. However, these systems target sub-6 GHz bands, where signals penetrate obstacles easily with minimal cell search overhead. At mmWave, car bodies block signals completely and narrow beams lose alignment easily, demanding fundamentally different solutions
Table 1. Comparison to existing works.
Related works Freq. (GHz) Mobility support Steer-able Tra/ Ref Soft HO Power (W) E2E impl.
[…]
Mobile mmWave Repeaters. Commercial mmWave repeaters for homes (Commware, 2020) and vehicles (Hyatt, [n. d.]) use external antennas to receive signals, amplify, and retransmit them to users inside. However, they consume 20-40 W, far exceeding Wall-Street’s 160 µW max power consumption [^2] With COTS DAC controllers (noa, [n. d.]), total power consumption ranges from 0.48 to 0.64 W. Custom DAC designs may further reduce the power consumption.. They cannot support concurrent transmission/reflection or make-before-break handover since they cannot relay two links at the same time. Such operation requires additional power-hungry RF chains. Moreover, repeaters requires higher search complexity compared to Wall-Street (Section˜4.2.1). Alternatively, rooftop-to-WiFi relay could bottleneck mmWave’s multi-Gbps to WiFi’s data rate while adding offloading latency.
Reconfigurable Intelligent Surface (RIS). RISs manipulate mmWave signals through programmable metamaterials that directly impart phase shifts on incoming waves rather than generating new signals, consuming minimal power. As seen in Table˜1, (Qian et al., 2022) is a passive RIS, reflecting signals to fixed directions for blind spots, unable to support mobile UEs. Other reflective RISs steer signals through switches (Tan et al., 2018), PIN diodes (Yang et al., 2022; Gros et al., 2021, 2023), or varactor diodes  (Wolff et al., 2023; Kim et al., 2024; Cho et al., 2023). While (Kim et al., 2024) integrates 5G protocols, it only reflects, not suitable for in-vehicle use. (Cho et al., 2023) is the first mmWave surface that can steer in both refraction and reflection but requires time-multiplexing between modes. Wall-Street supports simultaneous operation with arbitrary power division.
Roadside RIS deployments  (Chen et al., 2021; Alsenwi et al., 2022; Tian et al., 2022) could help gNBs overcome environmental blockage but require extensive infrastructure and cannot solve vehicle body penetration. Wall-Street, mounted directly on vehicles, moves with users to ensure continuous in-vehicle coverage and reliable HOs.
3. Primer: 5G SA Handover
Figure 2. 5G SA Xn handover: (1) Preparation phase; (2) Execution phase; (3) Completion phase. Data exchange is denoted with orange and yellow rectangles.
5G Standalone (SA) uses Xn handover, where gNBs exchange information related to the HO process via the Xn interface. The process has three phases as outlined in Fig.˜2: Preparation, Execution, and Completion.
(1) Preparation phase. 5G SA relies on UE-side feedback to trigger handovers. The serving cell periodically requests UEs to stop ongoing transmission and measure Reference Signal Received Power (RSRP) from neighboring cells. UEs search exhaustively, then send measurement reports (MR) containing neighbor IDs and RSRPs back to the serving gNB. This process disrupts service and drains UE battery.
A common trigger for 5G HO is event A3 (Deng et al., 2018; Xu et al., 2020). It triggers when neighbor cell RSRP ($M_{n}$) exceeds serving cell RSRP ($M_{s}$) by a certain hysteresis threshold $H$: $M_{n}>M_{s}+H$. To prevent ping-pong HO from signal fluctuations, HO only executes when A3 holds true for the time-to-trigger (TTT) duration.
(2) Execution phase. The serving gNB sends a handover request to the chosen target gNB via Xn interface. Upon acknowledgment, the serving gNB commands the UE to detach from itself and synchronize with the target using Random Access (RACH), while forwarding buffered data to the target gNB. However, handovers typically occur at edges where both links are weak, causing packet loss and retransmissions.
(3) Completion phase. The target gNB requests path switching to the 5G Core, then signals the source gNB to release the UE context via Xn interface. Poor HO decisions lead to ping-pong handovers or TCP connection termination.
Figure 3. (left) Wall-Street consists of over 4,000 paired electric and magnetic meta-atoms separated by a dielectric substrate. Biasing voltage on varactor diodes modulates current oscillation, creating EM field discontinuities; (right) VNA measured reflection and transmission response at 26 GHz, varying with biasing voltages ($u_{m}$, $u_{e}$).
4. Design
This section presents the design of Wall-Street, including the surface design (Section˜4.1) and HO system design (Section˜4.2). The serving gNB controls Wall-Street via sub-6 GHz control channel.
4.1. Wall-Street Hardware
Surface Design. Wall-Street is a 26 GHz Huygens metasurface (HMS) that supports both reflection and transmission. HMS consists of orthogonal electric and magnetic meta-atoms facing each other across a dielectric substrate (shown in Fig.˜3 on left), creating discontinuities in EM fields (Cho et al., 2023) [^3] We refer the reader to Appendix A for the details on our design parameters. Demonstration videos are available in bit.ly/42xb2nr. This structure can manipulate the magnitude and phase of the incident field in both reflective and transmissive directions. Each meta-atom contains a low-power voltage-controlled capacitor. By adjusting the bias voltages applied to the magnetic and electric meta-atoms $(u_{m},u_{e})$, we can change the complex reflection and transmission coefficients of the incoming signal. Figure˜3 (right) shows Vector Network Analyzer (VNA) measurements of Wall-Street’s response to different voltage combinations. The marked regions (black arrows) indicate where we achieve arbitrary phase shifts from 0 to $2\pi$ while maintaining high efficiency in reflection or transmission.
Dual-Link Beam Steering. Wall-Street’s key innovation is transflective operation; it simultaneously steers one beam in transmission (refraction) and another in reflection while adjustable power division between them. The challenge is to find an optimal configuration that jointly satisfies these multiple objectives, achieving two steering angles with a desired power ratio.
A straightforward approach is to spatially partition the surface into two - for example, assigning 80% of the area to the transmissive link and 20% to the reflective link. A distinct phase gradient is then applied to each partition to steer the two beams independently. While it is computationally fast, hard partitioning raises undesirable sidelobes and, based on our own experimental tests, reduces gain for both links.
Instead of hard partitioning, we use a soft partitioning approach that optimizes the contribution of individual meta-atom to both links. We use an optimization algorithm to find a single voltage configuration for the entire surface that maximizes the overall gain in the two desired directions while maintaining the power ratio and suppressing sidelobes. These configurations are pre-computed offline and stored in a codebook for real-time operation.
To populate this codebook, we find the optimal voltages $\left\{(u_{m,n},u_{e,n})\right\}_{n=0}^{N-1}$ for a desired pair of steering angles ($\theta_{1}$, $\theta_{2}$) with a power split ratio $\alpha$ using a genetic algorithm:
(1) $\max_{\{u_{m,n},u_{e,n}\}}~\Bigg|\sum_{n=0}^{N-1}\Big(\underset{\text{Transmissive link at }f_{1}}{\underbrace{\sqrt{1-\alpha}\,c_{t,n}\,e^{-j\phi_{t,n}}}}+\underset{\text{Reflective link at }f_{2}}{\underbrace{\sqrt{\alpha}\,c_{r,n}\,e^{-j\phi_{r,n}}}}\Big)\Bigg|^{2}$
where $\phi_{k}$ is the ideal phase shift required to steer a beam by $\theta_{k}$, and $c_{t,n}$, $c_{r,n}$ are the complex transmission and reflection coefficients of $n$-th meta-atom from S-parameter measurements (Fig.˜3). The resulting codebook $\Theta$, indexed by $(\theta_{t},\theta_{r},\alpha)$, is used during real-time operation (Section˜4.2). For dual-transmission, the reflective term is simply replaced with a second transmissive term, and for single-beam steering, one term is removed entirely.
(a) Cell Search (UL/DL)
(b) Make-Before-Break (UL/DL)
Figure 4. Wall-Street’s multi-beam operational modes: (a) dual-transflective steering for cell search and (b) dual-transmissive steering for make-before-break. Blue and red beams operate at different frequencies.
(a) Initial attachment.
(b) Handover Preparation: Measurement of candidate gNB 1 and gNB 2.
(c) Handover Execution: Make-before-break handover.
(d) Handover Completion.
Figure 5. System overview: (a) Initial UE-RAN attachment (Section˜4.2.1); (b) preparation using transflective surface power measurement (Section˜4.2.2); (c) beam combining to enable make-before-break dual-cell connectivity (Section˜4.2.4); (d) handover completion (Section˜4.2.5). Each gNBs use different frequencies, and the source gNB controls the surface.
Wall-Street’s dual-link beam steering, which independently controls two angles across a wide $\pm 70\degree$ range, provides two functions in our handover protocol, as seen in Fig.˜4. During neighbor cell search (Fig.˜4(a)), Wall-Street maintains a transmissive link between the serving gNB and UE (blue) while reflecting neighbor cell’s broadcast signals back toward the serving gNB (red) for cell search. During a make-before-break HO (Fig.˜4(b)), Wall-Street steers both the serving and target gNB’s signals to the UE. Angular reciprocity, where the surface maintains identical beam paths for uplink and downlink, enables instantaneous uplink/downlink switching without reconfiguration. In contrast, repeaters must switch between different receive and transmit beam patterns, adding latency.
(a) Initial attachment.
(b) Cell scanning.
Figure 6. Alignment process: initial attachment and cell scanning during HO preparation.
4.2. Wall-Street Handover Design
This section describes Wall-Street’s HO process. First, to ensure uninterrupted connectivity, we offload the neighbor cell measurement task from the UEs to the surface (Section˜4.2.2). Second, to improve efficiency, we batch the HO for all in-vehicle users into a single operation (Section˜4.2.3). Finally, to maximize reliability, we implement a make-before-break connection, linking the UE to the new cell before disconnecting from the old cell (Section˜4.2.4). The process is divided into: initial attachment, preparation, decision, execution, and completion, as detailed in Fig.˜5 (overview) and Fig.˜7 (timeline). Our design leverages the standard 5G practice where nearby gNBs transmit their broadcast signals on different time-frequency resources. This prevents collisions and allows the serving gNB to capture a neighbor’s reflected signal without interfering with the active data link.
4.2.1. Initial Attachment
As illustrated in Fig.˜6(a), the initial attachment requires a coordinated beam sweep to align the gNB, Wall-Street (operating in single-beam transmissive mode), and the UE. Every $20$ ms period, the gNB broadcasts a $5$ ms burst containing 64 synchronization signal blocks (SSBs), which are transmitted across 8 different gNB Tx beam directions. For each SSB, Wall-Street refract them into $8$ angles, testing different paths into the vehicle. For each burst, the UE holds its Rx beam and measures the power. The UE tests a different beam direction for each burst. After four bursts (one per UE Rx direction), taking $80$ ms total, the UE reports the best alignment and connects with the gNB.
A conventional amplify-and-forward repeater requires aligning four independent beams: the gNB Tx beam $f_{g}$, the repeater Rx beam $w_{r}$, the repeater Tx beam $f_{r}$, and the UE Rx beam $w_{u}$. This process can be written as:
(2) $\underset{\text{repeater (Tx beam $f_{r}$)}\to\text{UE}}{\underbrace{\left\|w_{u}^{H}\bm{H_{r\to u}}f_{r}\right\|^{2}}}\cdot\underset{\text{gNB}\to\text{repeater (Rx beam $w_{r}$)}}{\underbrace{\left\|w_{r}^{H}\bm{H_{g\to r}}f_{g}\right\|^{2}}}$
where $\bm{H_{r\to u}}$ and $\bm{H_{g\to r}}$ are channel blocks from repeater to UE and from gNB to repeater. Searching $(f_{g},w_{r},f_{r},w_{u})$ across four independent codebooks leads to $O(n^{4})$ complexity, taking about $\sim$5 sec, which is too slow for moving vehicles.
Wall-Street fundamentally reduces this complexity. It does not have separate Rx and Tx modules with independent beams. Instead, each surface configuration applies a specific phase gradient that deterministically maps any incoming wave to an outgoing wave (generalized Snell’s law: $k_{out}=k_{in}+\nabla_{\phi}(\theta)$). While the outgoing angle is dependent on the incoming angle for any single configuration, this does not constrain the search because the alignment process performs an full sweep of all gNB beam directions and all surface states. The received power is then formulated as:
(3) $\left\|w_{u}^{H}\bm{H_{s\to u}}\Theta(\theta)\bm{H_{g\to s}}f_{g}\right\|^{2}$
where $\bm{H_{s\to u}}$ and $\bm{H_{g\to s}}$ are channel blocks from surface to UE and from gNB to surface. Here, $\theta$ is a discrete surface state, and $\Theta$ is the Wall-Street’s codebook. This leads to three independent codebooks $(f_{g},\theta,w_{u})$ with $O(n^{3})$ complexity (Wu and Zhang, 2019; Jiang et al., 2024). Also, Wall-Street’s angular reciprocity allows fast UL/DL switching once a state is found.
Figure 7. Wall-Street’s handover timeline: (1) Preparation, (2) Execution, and (3) Completion phases.
4.2.2. Handover Preparation
The serving gNB periodically triggers Wall-Street to enter dual-beam mode for neighbor cell measurement. The goal is to measure neighbor cell RSRP without interrupting data sessions, as shown in Fig.˜5(b).
Since the refractive link between the serving gNB, Wall-Street, and the UE is already established, the transmissive angle $\theta_{t}$ is known. Wall-Street then uses its pre-computed codebook indexed by $(\theta_{t},\theta_{r},\alpha)$ to cycle through 8 reflective angles $\theta_{r}$, and neighbor gNBs transmit SSBs in their standard beam patterns. Since the serving gNB already knows its direction toward Wall-Street, it measures these reflected SSBs on a separate channel to prevent interference with the active data link. The power split $\alpha$ allocates less power to maintain the data connection while assigning the remaining power to the reflective link. As illustrated in Fig.˜6(b), this search takes one 5 ms burst. This allows the serving gNB to measure the neighbor RSRPs on behalf of the UEs without their involvement or service interruption.
4.2.3. Handover Decision.
After measuring neighbor cells, we determine if handover to the neighbor gNB improves signal quality for all UEs inside the vehicle. Standard A3 HO triggers when neighbor RSRP ($M_{n,i}$) exceeds serving RSRP ($M_{s,i}$) by a threshold $H$ ($M_{n,i}>M_{s,i}+H$) for $i$-th UE. However, this per-user HO could lead to conflicting decisions for different UEs. Also, the serving gNB measures the reflected neighbor signal ($M_{r}$), not the required $M_{n,i}$. As seen in Fig.˜8, we define the serving, neighbor, and reflected RSRP as:
(4) $M_{s,i}$ $=(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})+G_{\mathrm{w,t}}+(G_{ue,i}+L_{ue,i})$
(5) $M_{n,i}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,t}}+(G_{ue,i}+L_{ue,i})$
(6) $M_{r}$ $=(G_{\mathrm{gNB}_{n}}+L_{\mathrm{gNB}_{n}})+G_{\mathrm{w,r}}+(G_{\mathrm{gNB}_{s}}+L_{\mathrm{gNB}_{s}})$
The serving gNB directly measures $M_{s,i}$ from each UE and $M_{r}$ from Wall-Street. It also knows surface gains $G_{w,t}$ and $G_{w}$ from the codebook and the UE gain $G_{ue,i}$ [^4] In 3GPP, the gNB sends Downlink Control Information (DCI) to send power control commands to the UE.. The UE-specific path loss $L_{ue,i}$ remains unknown.
Instead of using gNB-UE links, our approach is to make one HO decision for the entire vehicle based on the predictable gNB-surface links. We redefine the serving link as $X_{s}=G_{gNB,s}+L_{gNB,s}$ and the neighbor link as $X_{n}=G_{gNB,n}+L_{gNB,n}$. Then, the HO decision for all $K$ users becomes $X_{n}>X_{s}+H$. Our goal is to estimate $X_{s}$ and $X_{n}$ from available measurements. To do so, we employ a two-slot measurement with bounding algorithm described next.
Figure 8. Measurement power translation: Inferring 3GPP-standard measurement report power readings based on measurements reflected from Wall-Street.
Two-slot measurements. We use Wall-Street’s dynamic power allocation to estimate $X_{s}$ and $X_{n}$ accurately. It allocates power between reflection ($\alpha_{j}$) and refraction ($1-\alpha_{j}$) in each time slot $j$. In Slot 1, we increase reflection to measure accurate $M_{r}$ (already measured in Section˜4.2.2), while in Slot 2, we increase refraction to accurately measure $M_{s,i}$. Based on Eq.˜4 and Eq.˜6, we define the per-slot models as:
(7) $M_{r}^{(j)}=X_{n}+X_{s}+G_{w,r}(\alpha_{j})$
(8) $M_{s,i}^{(j)}=X_{s}+G_{w,t}(1-\alpha_{j})+G_{ue,i}+L_{ue,i}^{(j)}$
Figure 9. Wall-Street-COSMOS testbed integration: we integrate Wall-Street with COSMOS USRP-based gNBs and UE (left). We illustrate a simplified control flow of Wall-Street-integrated COSMOS testbed (right).
Bounding algorithm. We use measurements $X_{s,i}$ from multiple UEs during Slot 2 to tightly bound $X_{s}$. We first subtract the known surface and UE gains from our measurements:
(9) $S_{1}=M_{r}^{(1)}-G_{w,r}(\alpha_{1})=X_{s}+X_{n}$
(10) $S_{2,i}=M_{s,i}^{(2)}-G_{w,t}(1-\alpha_{2})-G_{ue,i}=X_{s}+L_{ue,i}^{(2)}$
The path loss within the vehicle is physically bounded as $L_{min}\leq L_{ue,i}\leq L_{max}$. We can bound $X_{s}$ based on each UE measurement as $X_{s}\in[S_{2,i}-L_{max},S_{2,i}-L_{min}]$. By combining measurements from all UEs, we tighten this bound as:
$X_{s}\in[\underset{i\in K}{\text{max}}(S_{2,i}-L_{max}),\underset{i\in K}{\text{min}}(S_{2,i}-L_{min})]=[LB_{s},UB_{s}]$
Based on Eq.˜9, we can write our decision metric as $\Delta=X_{n}-X_{s}=S_{1}-2X_{s}$, so the bounds for our metric becomes $\Delta_{min}=S_{1}-2UBs$. We trigger HO if $\Delta_{min}\geq H$. Otherwise, HO does not occur to avoid unnecessary HOs. This conservative approach prevents rapid ping-pong HOs caused by transient signal fluctuations.
4.2.4. Handover Execution
The source gNB initiates the HO by sending a HO Req message to the target gNB. Upon receiving the HO ACK, the source gNB triggers Wall-Street to enter the make-before-break mode and sends HO command to the UE. The UE synchronizes with the target gNB via RACH and maintains two active protocol stacks using the 3GPP-standardized Dual Active Protocol Stack (DAPS) (3GPP, 2021). It establishes new Signaling Radio Bearers (SRBs) for control-plane signaling (including Wall-Street control) with the target cell while suspending the source cell SRBs. For the user plane, a common Packet Data Convergence Protocol (PDCP) entity is created to serve both links, allowing the UE to exchange data with both gNBs on different channels.
At the physical layer, Wall-Street combines downlink beams from both gNBs and splits the uplink beam towards both gNB. No additional beam searching is required for this step as the optimal angles for the serving link ($\theta_{1}$) and the target link ($\theta_{2}$) are already known from HO preparation. To ensure reliable packet delivery, both the source and target gNBs transmit duplicate data packets to the UE, where the PDCP layer handles reordering and de-duplication.
Wall-Street initially divides power equally between the two links. If one link degrades, it temporarily boosts that path to maintain a minimum quality of service, otherwise it falls back to an even split. During make-before-break, Wall-Street can directly measure both $M_{s}$ and $M_{n}$ to check A3 condition in real time. If the condition is not met for a TTT duration, the system reverts to the serving cell.
4.2.5. Handover Completion
The target cell completes the HO by sending a UE release to the source cell and HO completion to Wall-Street. Upon completion, Wall-Street maximizes transmission power to the new gNB. If the HO fails, Wall-Street instead maximizes power to the old gNB so that UE does not restart from the initial attachment phase. The UE also seamlessly reverts to the source gNB configuration and reactivates source cell SRBs for control-plane signaling. We note that a single HO execution simultaneously transfers all UEs within the same vehicle, reducing control-plane signaling.
(a) Wall-Street hardware
(b) Indoor setup
(c) Outdoor in-vehicle multi-UE setup
Figure 10. Wall-Street’s hardware implementation and testbed: showing (a) individual hardware components, (b) indoor experimental UE-side setup, and (c) outdoor in-vehicle UE-side configuration.
(a) Indoor setup
(b) Outdoor setup
Figure 11. Experimental gNB-side testbed setup and map: we deploy gNBs on the first floor of a lab facing an outdoor parking lot. The vehicle is approximately 2020 to $30$ m from the gNBs. The arrow indicates the direction that each gNB node is facing at 0°, and each node sweeps the beam from -60° to 60°.
5. Implementation
Figure˜9 illustrates our implementation of Wall-Street. It consists of (1) a surface hardware and control unit; (2) integration with the COSMOS testbed, implementing new PHY features; and (3) an ns-3 simulation for evaluating TCP end-to-end performance, using a state-of-the-art mmWave ns-3 simulator (Mezzavilla et al., 2018) with our own modifications. The simulation is trace-driven, fed with PHY traces collected from COSMOS testbed in real-time under a multi-user scenario.
5.1. Wall-Street Hardware
The Wall-Street surface is composed of 76 boards, each consisting of 28 co-located magnetic and electric unit cells as shown in Fig.˜10(a). Two 40-channel AD5370 16-bit DACs provide independent control of the unit cells on each Wall-Street board, with each channel supplying a variable 0 to 16 V control voltage. To expedite the steering process, one channel supplies the voltage to two adjacent boards, enabling two DACs to independently control all boards. The two DACs are each connected to the Serial Peripheral Interface (SPI) of a Raspberry Pi (RPI) through GPIO. The RPI listens for control signals from the COSMOS testbed control network, known as sandbox1 (SB1), via sub-6 GHz. Upon receiving a signal, it sends a command to the DACs, which then apply the appropriate voltage levels. These voltage levels are determined from a pre-stored angle-to-voltage codebook. For multi-beams, we construct codebook for different power ratios between two beams ($\alpha=0.25,0.5,0.75$). The speed of the Wall-Street hardware has been optimized to 0.2 ms for real-time experiments.
5.2. COSMOS Testbed
The COSMOS testbed comprises software-defined radios (USRPs) integrated with the IBM mmWave Phased-Array Antenna Module (PAAM) frontends (Fig.˜9). The PAAM features 4 tiled RFICs, each with 32 TRx phase-shifting elements, supporting a total of 64 antennas. The PAAM offers full access to beamforming control and latency as low as 10s of nanoseconds for low-latency MAC and hybrid beamforming. As shown in Fig.˜10(a), our experimental setup uses two types of PAAM devices: two stationary PAAM nodes integrated with USRP N310s mounted on remotely controllable XY tables, and three portable PAAM nodes integrated with USRP 2974s. For indoor scenarios, we utilize four gNBs: two stationary PAAMs (gNB 1 and 4) with a 35 dB gain and two portable PAAMs (gNB 2 and 3) with a 31 dB gain. In multi-UE outdoor scenarios, we use three gNBs: two stationary PAAMs (gNB 1 and 2) with a 65 dB gain and one portable PAAM (gNB 3) with a 31 dB gain, with the remaining portable PAAMs serving as UEs with a 31 dB gain [^5] The node gains are defined by the COSMOS testbed. We refer the reader to https://www.cosmos-lab.org/technology/mmwave/ for the details.. The location of UEs remain the same with and without the surface. Since all nodes must be connected to the same control network (SB1), we use a Wi-Fi bridge operating at sub-6 GHz to send the control commands to outdoor UEs and the surface.
5.2.1. PHY implementation on COSMOS
We developed new features on the PHY layer of the COSMOS testbed, including OFDM packet transmission and reception on two different channels, extraction of RSRP and Packet Error Rate (PER) from the packet, and IBM PAAM and Wall-Street beam control. Periodic bursts of OFDM packets are sent from two PAAMs in different frequency bands with 3.84 MHz separation, and the receiver simultaneously decodes packets from both transmitters by operating at a larger bandwidth. Specifically, the receiver extracts RSRP from the reference signal in the preamble, the packet sequence number from the header, and the PER from the cyclic redundancy check (CRC). During the make-before-break handover, the source and target gNBs transmit identical packets, and the UE decodes packets on both channels simultaneously and compute the combined PER. A C++ script in SB1 sends control commands to the gNB and Wall-Street to follow the handover procedure, as illustrated in Fig.˜9 (right).
Figure 12. Beam patterns for four steering modes. Red/blue $\triangle$ show measurements at 26.0/26.1 GHz. $\blacksquare$ show baseline (surface turned off). Gray lines indicate target angles. Left: -40°/40° steering; right: -45°/68° steering.
6. Evaluation
We first present microbenchmark on beam pattern, coverage improvement, and dual-link capabilities of Wall-Street’s (Section˜6.2). We then conduct field studies to evaluate the end-to-end performance in mobile scenarios, comparing Wall-Street against standard 5G SA handover protocol (Section˜6.3).
6.1. Methodology
We evaluate Wall-Street in both controlled indoor (single-UE) and real-world outdoor (multi-UE) driving scenarios. For indoor settings, there are four gNB nodes and one UE node in a $20\times 30$ m lab, as shown in Fig.˜11(a). To test the UE and Wall-Street under mobility, we place them on a cart moving in a constant speed of approximately 1 km/h. For outdoor settings, there are three gNB nodes on the first floor of a lab facing a parking lot. We mount the surface on the rear window of a SUV vehicle and deploy two UE nodes. UE 1 is on the rear seat (opposite side from surface), and UE 2 is on the cargo area, as shown in Fig.˜10. The vehicle drives at speeds ranging from 5 to 15 km/h. Both the surface and UEs are battery-powered, with UE positions remaining the same for the baseline. Since the gNB nodes use stationary PAAMs that cannot be relocated, our outdoor driving trajectory is limited to 30 m, as depicted in Fig.˜11.
We implement the 5G SA HO protocol as our baseline and use this as a performance benchmark. For both Wall-Street and baseline, the hysteresis parameter $H$ of A3 event is set to $10$ dB with TTT of $150$ ms and the measurement report periodicity of $160$ ms. For the baseline, a total duration of every measurement report takes less than $20$ ms with three neighbor gNBs. Also, we set the maximum in-vehicle distance to $1$ m for $L_{min}$. Our results combine (1) PHY measurements, including RSRP and packet error rate, from COSMOS testbed with real vehicle trajectory, and (2) TCP evaluation using ns-3 simulator driven by PHY traces. In the ns-3 simulator, we increased all outdoor RSRP traces by 10 dB for both the baseline and Wall-Street scenarios.
6.2. Microbenchmarks
We evaluate Wall-Street’s far-field dual-beam pattern measured with the spectrum analyzer, in-vehicle signal coverage, and dual-beam operations at the physical layer.
Far-field beam patterns. Figure˜12 demonstrates Wall-Street’s beam steering capabilities in four configurations: dual-transmissive mode where both beams pass through the surface (Fig.˜12 (a-b)), and transflective mode where one beam transmits through while another reflects (Fig.˜12 (c-d)). Each configuration shows both uplink and downlink operations, validating bidirectional functionality.
We evaluate two beam angle pairs (-40°/40°, -45°/68°), representing typical handover scenario (see Appendix˜B for angle selections). To measure beam patterns, we place Tx and Rx horn antennas two meters from Wall-Street and sweep the Rx from -70°to 70°. Two transmitters operate on separate frequencies, 26.0 GHz (blue $\triangle$) and 26.1 GHz (red $\triangle$), enabling simultaneous beam measurement. Gray lines mark the target steering angles. Without an anechoic chamber, a large portion of signal power diffracts around the surface. We quantify this baseline leakage by disabling Wall-Street (0V, shown as $\blacksquare$) and normalize all measurements to the spectrum analyzer’s noise floor (0 dB reference).
Figure 13. Power allocation between beams at -40°/55°, normalized to baseline. Balanced (50/50) vs. asymmetric (75/25) distribution. Gray lines indicate target angles.
Results show 20-25 dB gain over baseline at target angles, except when a steered beam coincides with the natural specular reflection angle (visible at -40°/40° in Fig.˜12(c) and Fig.˜12(d)). Gain decreases at wider steering angles. We use identical voltage configurations for both uplink and downlink (e.g., Fig.˜12(a) and Fig.˜12(b)), confirming angular reciprocity that allows fast UL/DL switching during handover.
Beam power ratio. Figure˜13 shows dynamic power allocation between beams at -40°/55°. All measurements are normalized against the baseline (surface off) for visibility. Shifting from balanced (50/50) to asymmetric (75/25) distribution increases the first beam (blue) while decreasing the second beam (red) at target angles. Both beams maintain target angles (gray lines) and only relative power changes.
In-vehicle mmWave coverage. We evaluate Wall-Street’s ability to overcome vehicle body blockage by mounting it on a vehicle’s rear window with a gNB positioned 30 m away. For each in-vehicle location, we measure maximum achievable RSRP by exhaustively searching all beam combinations. In Figure˜14, nearly 50% of the vehicle interior experiences complete signal outage without Wall-Street. On the other hand, Wall-Street ensures outage-free coverage inside the vehicle and provides signal strength improvement of over 12 dB.
Figure 14. In-vehicle mmWave coverage. $\bullet$: no signal.
(a) Outdoor coverage for UE 1
(b) Outdoor coverage for UE 2
Figure 15. Signal strength and beam angles during vehicle movement. Upper: RSRP; middle: corresponding gNB beam angles; lower: UE/surface angles.
Beam tracking under mobility. Figure˜15 shows RSRPs and corresponding beam angles for gNB and UE/surface as the vehicle moves along a 30-meter outdoor path for UE 1 (Fig.˜15(a)) and UE 2 (Fig.˜15(b)). Wall-Street provides significantly higher signal strength overall. This improvement becomes particularly evident for UE 2. Without Wall-Street, UE 2 loses signals 70-80% of the time despite three available gNBs, whereas with Wall-Street, it receives consistent power from either gNB 1 or gNB 3 throughout the entire path.
(a) Baseline single-link
(b) Wall-Street dual-links
Figure 16. Indoor multi-link evaluation: (a) baseline single-link: optimal beam alignment (upper), misaligned (lower). (b) Wall-Street dual-link: make-before-break with 50/50 power split (upper), RSRP inference from reflected signals (lower). gNB $\blacktriangle$, blockage $\blacksquare$.
Multi-link operations. Figure˜16 evaluates Wall-Street’s ability to handle multiple simultaneous beams at various indoor locations. Figure˜16(a) shows baseline single-link performance without Wall-Street. The upper heatmaps demonstrate best-case RSRP when UE beams are optimally aligned with each gNB. The lower heatmaps show worst-case RSRP when beam is misaligned. Figure˜16(b) demonstrates Wall-Street’s dual-link capabilities. The upper heatmaps show dual-link signal reception where Wall-Street splits power equally between gNB1 and gNB4. While individual RSRPs are about 3dB lower than optimal alignment due to power division, both links remain active. The lower heatmaps validate our RSRP inference method (Section˜4.2.3) by showing signal strengths calculated from reflected measurements.
(a) Indoor TCP results with one UE.
(b) Outdoor TCP results with two UEs (upper: UE 1; lower: UE 2).
Figure 17. Throughput and round trip time (RTT) achieved by four congestion control algorithms. The right and lower edge of the box represents the 10% percentile, respectively. The left and upper edge give the 90% percentiles. The intersection point of the horizontal and vertical error bar represents the median value.
6.3. End-to-end Throughput and Delay
We evaluate end-to-end performance of Wall-Street delivering bulk TCP data flows. To quantify performance, we calculate throughput as the total data bits divided by the duration, measured every 100 ms. We average the round-trip-time (RTT) of packets over every 100 ms.
TCP performance under various settings. We evaluate the throughput and RTT of Wall-Street and the baseline under indoor and outdoor settings using four congestion control algorithms: CUBIC, Vegas, HighSpeed, and NewReno. In the indoor setting (Fig.˜17(a)), Wall-Street enhances average throughput by at least 20% and reduces delay by 20-80% compared to the baseline. For the outdoor multi-user scenario (Fig.˜17(b)), Wall-Street yields more significant improvements. Throughput improves by up to 78% (68% on average for UE 1 and 60% on average for UE 2) and RTT reduction by up to 34% (14% on average for UE 1 and 5% on average for UE 2). Notably, in the baseline setup, UE 2 (located in the cargo area) experiences lower throughput and higher RTT than UE 1 due to car body blockage.
TCP performance over time. Figure˜18 divides the 20-second outdoor trajectory into 10 intervals, showing median throughput and delay for UE 1 (upper) and UE 2 (lower). Vertical dashed lines mark handover events. Wall-Street demonstrates three key advantages over baseline. First, Wall-Street maintains higher throughput and lower latency throughout the trajectory, with UE 1 achieving higher throughput improvement and UE 2 maintaining connectivity where baseline fails. Second, at the 6s mark, Wall-Street executes a single batch handover for both UEs simultaneously. This contrasts with baseline, where UE 1 and UE 2 trigger separate handovers at different times, causing individual service interruptions and increased signaling overhead. Lastly, baseline suffers from ping-pong handovers between gNB 1 and gNB 3 and a complete connection loss for UE 2 at 12s following a failed handover. Wall-Street eliminates both issues through stable decision-making and make-before-break execution.
Figure 18. Throughput and delay of UE 1 (upper) and UE 2 (lower) achieved without Wall-Street (left) and with Wall-Street (right) as the vehicle is moving along the same trajectory. Vertical dashed lines indicate the handover event.
Figure 19. Impact of speed from 5 km/h to 15 km/h: throughput, RTT, HO count, and total outage duration (left to right).
Impact of vehicle speed. Figure˜19 compares Wall-Street’s performance against the baseline at driving speeds from 5 km/h to 15 km/h. Wall-Street consistently outperforms the baseline across most metrics and speed scenarios. It improves user throughput by 70% at 5 km/h, 27% at 10 km/h, and 8.4% at 15 km/h. It also makes the network connection more reliable by reducing the number of handovers by an average of 37% and delay by an average of 11% and cutting connection outages by an average of 64%. Wall-Street maintains the same low number of handovers regardless of speed because it makes one conservative decision for all users simultaneously based on multiple measurements, preventing the unnecessary handovers that happens the baseline. However, performance gains diminish at higher speeds due to a limitation of our current implementation where the system cannot complete handover preparation quickly enough. We discuss potential solutions to accelerate the scanning in Section˜7.
Impact of handover preparation. Figure˜20 illustrates sequence number progression during the outdoor experiment, with the left subfigure showing the entire duration and the right subfigure focusing on the handover preparation phase. To enhance visibility of trends, sequence numbers are subjected to a modulo operation. Wall-Street allows a steeper increase in sequence numbers compared to the baseline, indicating that more packets are successfully delivered. During handover preparation, the baseline’s sequence numbers show irregular progression and a lower slope due to service interruptions as the UE stops communicating with the source gNB to detect SSB bursts from neighboring gNBs. While SSB detection takes under 20 ms, its effects persist for 50 ms, significantly disrupting baseline packet delivery. In contrast, Wall-Street maintains steady sequence number progression throughout the handover preparation phase, demonstrating more stable communication.
Figure 20. Measurement report impact: increase trend of sequence numbers during the entire 30s experiment (left) and handover preparation (right).
Impact of make-before-break handover. Figure˜21 evaluates make-before-break performance using real-time packet error rate (PER) measurements during a 25-second drive at 5 km/h. Both gNB1 and gNB3 transmit duplicate packets (identified by sequence numbers), enabling per-link and combined PER calculation over time.
The baseline setup faces a fundamental limitation where UE 1 can only align with one gNB despite both transmitting. We configure UE 1’s beam toward gNB 3 while both gNBs beamform toward UE 1. This single-alignment constraint means gNB 1’s packets rarely reach the UE, negating make-before-break benefits.
Wall-Street eliminates this limitation by simultaneously steering both gNB signals to the UE. The upper plots show individual link PER with gNB1 (red) and gNB3 (blue), while lower plots show combined PER when receiving packets from both source (black).
Without Wall-Street, gNB 3 provides relatively good PER but gNB 1 remains mostly unusable, rarely benefiting from make-before-break mechanism. With Wall-Street, both links operate at moderate PER due to power splitting, but their combination achieves low PER throughout the trajectory. This validates that Wall-Street transforms make-before-break from a theoretical capability into practical reliability.
Figure 21. Make-before-break impact: real-time packet error rates (PERs) of a link from each gNB (upper) and that of two links with duplicated packets from both gNBs (lower).
7. Discussion
In this section, we discuss Wall-Street’s current limitations and potential solutions.
Control Channel. Our prototype uses an out-of-band sub-6 GHz link to integrate with the COSMOS roadside testbed, as the testbed requires all components to be connected through the same local network. However, Wall-Street only requires a few control bits to index its pre-computed voltage codebook. In deployment, this control information can move in-band. The gNB can embed low-rate control data directly into the mmWave Synchronization Signal Blocks (SSBs) it already transmits, using simple modulation schemes such as Frequency-Shift Keying (FSK) on a subcarrier. The surface would then need only a low-power detector synchronized to the SSB periodicity to decode these commands, eliminating the need for an active modem entirely.
Reducing Alignment Overhead. RIS-assisted roadside networking faces significant beam alignment overhead across the gNB, surface, and UE. We can reduce this complexity through two approaches. First, vehicles’ existing GPS systems provide rough location estimates that allow the gNB to narrow its transmit beam search space, effectively eliminating one search dimension during the beam alignment. We note that we may use GPS only for accelerating beam alignment, but not for handover decisions as location alone cannot capture environmental blockages or vehicle body effects that determine actual signal quality between the base station and UEs. Second, the surface can operate as a reflective probe during the initial alignment. It can reflect and steer the gNB’s SSB transmissions while the gNB listens for the strongest return, taking $O(N^{2})$ to determine their mutual orientation. The surface can then reflect and steer UE transmission toward the UE with another $O(N^{2})$ search. This two-stage approach could reduce the complexity of initial alignment.
8. Conclusion
This paper introduces the first smart surface-enabled 5G mmWave roadside networking system designed to efficiently steer outdoor mmWave signals into vehicles, enable simultaneous data exchange and neighboring cell measurements and reliable make-before-break handovers. We conduct an extensive evaluation in various indoor and outdoor multi-user settings, demonstrating significant TCP improvement. Future work will explore Wall-Street’s performance under various environmental conditions and mobility scenarios, including different weather and terrain variations.
Appendix A Surface Design Parameters
In this section, we provide design parameters of Wall-Street’s electric and magnetic meta-atom, as illustrated in Fig.˜22. The thickness of PCB Roger 4003C substrate is $0.406$ mm and the thickness of the copper trace is $1$ oz. To simulate the unit cell response, we create an air box that defines the boundary conditions. The top and bottom boundaries of the box are perfect electric boundary, and the front and back of the box are perfect magnetic boundary.
Appendix B Beam Angle Selection
We conduct Wall-Street beam pattern microbenchmarks using angle pairs of -40°/40°, -40°/55°, and -68°/45°. These angles reflect typical mmWave handover scenarios, where handovers occur every 0.13 km within a gNB’s typical coverage radius of 0.15 km (Hassan et al., 2022). The demonstration video of different multi-beam angles are also available in bit.ly/42xb2nr.
Figure 23. Beam angle selection for microbenchmark.